# An Inexact Bregman Proximal Difference-of-Convex Algorithm with Two Types of Relative Stopping Criteria

### v1

@@ 5.2 The constrained $\ell_{1-2}$ sparse optimization problem
where $\mu\in[0,1)$, $A\in\mathbb{R}^{m\times n}$ has full row rank, and $\bm{b}\in\mathbb{R}^{m}$. Moreover, we choose $\kappa\in(0,\,\|\bm{b}\|)$ so that the origin is not a feasible point and follow [55, Section 6] to choose $M:=(1-\mu)^{-1}\left(\|A^{{\dagger}}\bm{b}\|_{1}-\mu\|A^{\dagger}\bm{b}\|\right)$ so that the feasible region is bounded, where $A^{{\dagger}}$ denotes the pseudo-inverse of $A$. To apply our iBPDCA in Algorithm 1 for solving problem (5.8), we consider the following equivalent reformulation:
@@
$\min_{\bm{x}\in\mathbb{R}^{n}}~{}~{}\iota_{\kappa}(A\bm{x}-\bm{b})+\iota_{M}(\bm{x})+\|\bm{x}\|_{1}-\langle\boldsymbol{\xi}^{k},\,\bm{x}-\bm{x}^{k}\rangle+\frac{\gamma_{k}}{2}\big{\|}\bm{x}-\bm{x}^{k}\big{\|}^{2}+\frac{\gamma_{k}}{2}\big{\|}A\bm{x}-A\bm{x}^{k}\big{\|}^{2},$ (5.9)
@@
$\min_{\bm{z}\in\mathbb{R}^{m}}\left\{\begin{aligned} \Psi_{k}^{\text{con}}(\bm{z})&:=\langle\boldsymbol{z},\bm{b}\rangle+\frac{\gamma_{k}}{2}\left\|\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right\|^{2}-\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|_{1}\\[3.0pt] &\qquad-\frac{\gamma_{k}}{2}\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)-\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|^{2}\\[3.0pt] &\qquad+\frac{\gamma_{k}}{2}\left\|\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right\|^{2}-\frac{\gamma_{k}}{2}\left\|\Pi_{\kappa}\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)-\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)\right\|^{2}\\[3.0pt] &\qquad-\frac{\gamma_{k}}{2}\|\bm{s}^{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bm{b}^{k}\|^{2}\end{aligned}\right\},$ (5.11)
where $\bm{z}\in\mathbb{R}^{m}$ is the dual variable and $\Pi_{\kappa}$ is the projection operator over $\left\{\bm{x}\in\mathbb{R}^{n}:\|\bm{x}\|\leq\kappa\right\}$. From the property of the Moreau envelope of $\gamma_{k}^{-1}g$ and $\gamma_{k}^{-1}\iota_{\kappa}$ (see, e.g., [6, Proposition 12.29]), we see that $\Psi_{k}^{\text{con}}$ is convex and continuously differentiable with the gradient:
@@
and $\partial\Pi_{\kappa}\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)$ is the Clarke subdifferential of the Lipschitz continuous mapping $\Pi_{\kappa}(\cdot)$ at $\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}$, defined as follows (see, e.g., [58, Remark 3.1]):
@@
Note that the elements in ^$\widehat{\partial}^{2}\Psi_{k}^{\text{con}}(\bm{z})$ may only be positive semidefinite. We then need to incorporate a small adaptive regularization term when solving the linear system in the Ssn method. We present the whole iterative framework in Algorithm 3 and refer readers to [59, Theorems 3.4 and 3.5] for its convergence results.
@@
We next show that our inexact stopping criteria (SC1) and (SC2) can also be achieved through the appropriate manipulations based on the dual sequence generated by the Ssn method in Algorithm 3. To this end, we first assume that a partially strictly feasible point $\bm{x}^{\text{feas}}$ satisfying $\|A\bm{x}^{\text{feas}}-\bm{b}\|<\kappa$ and $\|\bm{x}^{\text{feas}}\|_{\infty}\leq M$ is available on hand. Indeed, from [54, Section 6.1], with the choice of $M=(1-\mu)^{-1}\left(\|A^{{\dagger}}\bm{b}\|_{1}-\mu\|A^{\dagger}\bm{b}\|\right)$, such a point can be simply obtained by setting $\bm{x}^{\text{feas}}:=A^{\dagger}\bm{b}$. Then, at the $k$-th iteration, we apply the Ssn method for solving equation (5.12), which generates a dual sequence $\{\bm{z}^{k,t}\}$. Let
@@
Similar to discussions at the end of subsection 5.1.1, one can also see from Proposition 5.2 that our inexact stopping criterion (SC1) or (SC2) holds as long as the checkable inequality (5.13) or (5.14) holds. Moreover, when the dual sequence $\{\bm{z}^{k,t}\}$ generated by the Ssn method is convergent under proper conditions (see [59, Theorems 3.4 and 3.5]), we have that $\{\bm{w}^{k,t}\}$ is bounded, $\bm{e}^{k,t}\to 0$ and $\widetilde{\bm{w}}^{k,t}-\bm{w}^{k,t}\to 0$. Thus, we see that $\|\Delta^{k,t}\|^{2}+|\langle\Delta^{k,t},\,\widetilde{\bm{w}}^{k,t}-\bm{x}^{k}\rangle|+\delta^{1}_{k,t}+\delta^{2}_{k,t}\to 0$. On the other hand, when $\bm{x}^{k}$ is not the optimal solution of the $k$-th subproblem (5.9) or $\bm{x}^{k}\neq\bm{x}^{k-1}$, the right-hand-side term in (5.13) or (5.14) cannot approach zero. Therefore, the inequality (5.13) or (5.14) must hold after finitely many iterations.
@@ 5.2.2 Comparison results
We follow [55, Section 6.1] to choose the extrapolation parameters $\{\beta_{k}\}$ and other involved parameters.
@@
Finally, we test all the methods with two instances $(A,\bm{b})$ obtained from the data sets mpg7 and gisette in the UCI data repository. For each data set, we choose $\kappa=\kappa_{c}\|\bm{b}\|$ with $\kappa_{c}\in\{0.1,0.2\}$. The numerical results are reported in Figure 2, where the normalized objective value is computed by $\frac{F_{\text{con}}(\bm{x}^{k})-F_{\text{con}}^{\min}}{F_{\text{con}}(\bm{x}^{0})-F_{\text{con}}^{\min}}$ with $F_{\text{con}}^{\min}$ being the minimum of the terminating objective values obtained among all methods. The results further demonstrate the encouraging performance of our iBPDCA-SC1 and iBPDCA-SC2 on the real data sets.
@@ 6 Conclusions
Appendix A Missing proofs in Section 5
A.1 Proof of Proposition 5.1

### v2

@@ 5.2 The constrained $\ell_{1-2}$ sparse optimization problem
where $\mu\in[0,1)$, $A\in\mathbb{R}^{m\times n}$ has full row rank, and $\bm{b}\in\mathbb{R}^{m}$. Moreover, we choose $\kappa\in(0,\,\|\bm{b}\|)$ so that the origin is not a feasible point and follow [58, Section 6] to choose $M:=(1-\mu)^{-1}\left(\|A^{{\dagger}}\bm{b}\|_{1}-\mu\|A^{\dagger}\bm{b}\|\right)$ so that the feasible region is bounded, where $A^{{\dagger}}$ denotes the pseudo-inverse of $A$. To apply our iBPDCA in Algorithm 1 for solving problem (5.8), we consider the following equivalent reformulation:
@@
$\min_{\bm{x}\in\mathbb{R}^{n}}~{}~{}\iota_{\kappa}(A\bm{x}-\bm{b})+\iota_{M}(\bm{x})+\|\bm{x}\|_{1}-\langle\bm{\xi}^{k},\,\bm{x}-\bm{x}^{k}\rangle+\frac{\gamma_{k}}{2}\big{\|}\bm{x}-\bm{x}^{k}\big{\|}^{2}+\frac{\gamma_{k}}{2}\big{\|}A\bm{x}-A\bm{x}^{k}\big{\|}^{2},$ (5.9)
@@
$\min_{\bm{z}\in\mathbb{R}^{m}}\left\{\begin{aligned} \Psi_{k}^{\text{con}}(\bm{z})&:=\langle\bm{z},\bm{b}\rangle+\frac{\gamma_{k}}{2}\left\|\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right\|^{2}-\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|_{1}\\[3.0pt] &\qquad-\frac{\gamma_{k}}{2}\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)-\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|^{2}\\[3.0pt] &\qquad+\frac{\gamma_{k}}{2}\left\|\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right\|^{2}-\frac{\gamma_{k}}{2}\left\|\Pi_{\kappa}\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)-\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)\right\|^{2}\\[3.0pt] &\qquad-\frac{\gamma_{k}}{2}\|\bm{s}^{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bm{b}^{k}\|^{2}\end{aligned}\right\},$ (5.11)
where $\bm{z}\in\mathbb{R}^{m}$ is the dual variable and $\Pi_{\kappa}$ is the projection operator over $\left\{\bm{x}\in\mathbb{R}^{n}:\|\bm{x}\|\leq\kappa\right\}$. The detailed derivation of the dual problem (5.11) is relegated to Appendix A.2. From the property of the Moreau envelope of $\gamma_{k}^{-1}g$ and $\gamma_{k}^{-1}\iota_{\kappa}$ (see, e.g., [6, Proposition 12.29]), we see that $\Psi_{k}^{\text{con}}$ is convex and continuously differentiable with the gradient:
@@
and $\partial\Pi_{\kappa}\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)$ is the Clarke subdifferential of the Lipschitz continuous mapping $\Pi_{\kappa}(\cdot)$ at $\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}$, defined as follows (see, e.g., [61, Remark 3.1]):
@@
Note that the elements in ^$\widehat{\partial}^{2}\Psi_{k}^{\text{con}}(\bm{z})$ may only be positive semidefinite. We then need to incorporate a small adaptive regularization term when solving the linear system in the Ssn method. We present the whole iterative framework in Algorithm 3 and refer readers to [62, Theorems 3.4 and 3.5] for its convergence results.
@@
We next show that our inexact stopping criteria (SC1) and (SC2) can also be achieved through the appropriate manipulations based on the dual sequence generated by the Ssn method in Algorithm 3. To this end, we first assume that a partially strictly feasible point $\bm{x}^{\text{feas}}$ satisfying $\|A\bm{x}^{\text{feas}}-\bm{b}\|<\kappa$ and $\|\bm{x}^{\text{feas}}\|_{\infty}\leq M$ is available on hand. Indeed, from [57, Section 6.1], with the choice of $M=(1-\mu)^{-1}\left(\|A^{{\dagger}}\bm{b}\|_{1}-\mu\|A^{\dagger}\bm{b}\|\right)$, such a point can be simply obtained by setting $\bm{x}^{\text{feas}}:=A^{\dagger}\bm{b}$. Then, at the $k$-th iteration, we apply the Ssn method for solving equation (5.12), which generates a dual sequence $\{\bm{z}^{k,t}\}$. Let
@@
Similar to discussions at the end of subsection 5.1.1, one can also see from Proposition 5.2 that our inexact stopping criterion (SC1) or (SC2) holds as long as the checkable inequality (5.13) or (5.14) holds. Moreover, when the dual sequence $\{\bm{z}^{k,t}\}$ generated by the Ssn method is convergent under proper conditions (see [62, Theorems 3.4 and 3.5]), we have that $\{\bm{w}^{k,t}\}$ is bounded, $\bm{e}^{k,t}\to 0$ and $\widetilde{\bm{w}}^{k,t}-\bm{w}^{k,t}\to 0$. Thus, we see that $\|\Delta^{k,t}\|^{2}+|\langle\Delta^{k,t},\,\widetilde{\bm{w}}^{k,t}-\bm{x}^{k}\rangle|+\delta^{1}_{k,t}+\delta^{2}_{k,t}\to 0$. On the other hand, when $\bm{x}^{k}$ is not the optimal solution of the $k$-th subproblem (5.9) or $\bm{x}^{k}\neq\bm{x}^{k-1}$, the right-hand-side term in (5.13) or (5.14) cannot approach zero. Therefore, the inequality (5.13) or (5.14) must hold after finitely many iterations.
@@ 5.2.2 Comparison results
We follow [58, Section 6.1] to choose the extrapolation parameters $\{\beta_{k}\}$ and other involved parameters.
@@
Finally, we test all the methods with two instances $(A,\bm{b})$ obtained from the data sets mpg7 and gisette in the UCI data repository. For each data set, we choose $\kappa=\kappa_{c}\|\bm{b}\|$ with $\kappa_{c}\in\{0.05,0.1,0.2\}$. The numerical results are reported in Figure 2, where we plot the objective value $F_{\text{con}}(\bm{x}^{k})$ against the computational time. The results further demonstrate the encouraging performance of our iBPDCA-SC1 and iBPDCA-SC2 on real data sets.
@@ 6 Conclusions
Note that, in [45], the authors also proposed a Bregman proximal DC algorithm with extrapolation (BPDCAe), which demonstrated better numerical performance compared to BPDCA. Exploring the potential extension of the BPDCAe algorithm to an inexact version is both interesting and valuable. However, such an extension would require substantial additional effort and is beyond the scope of the current paper. We will consider this as a future research topic.
Acknowledgments
We thank the editor and referees for their valuable suggestions and comments, which have helped to improve the quality of this paper.
Appendix A Missing proofs and derivations in Section 5
A.1 Derivation of the dual problem (5.4)
First, the associated Lagrangian function of (5.3) is given by
$\mathcal{L}(\bm{x},\bm{y},\bm{z})$ $=\lambda\|\bm{x}\|_{1}-\langle\bm{\xi}^{k},\bm{x}\rangle+\frac{1}{2}\|\bm{y}-\bm{b}\|^{2}+\frac{\gamma_{k}}{2}\|\bm{x}-\bm{x}^{k}\|^{2}+\langle\bm{z},\bm{\bm{A}}\bm{x}-\bm{y}\rangle$
$=\lambda\|\bm{x}\|_{1}+\frac{\gamma_{k}}{2}\|\bm{x}-\bm{v}_{k}(\bm{z})\|^{2}-\frac{\gamma_{k}}{2}\left\|\bm{v}_{k}(\bm{z})\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{x}^{k}\|^{2}+\frac{1}{2}\|\bm{y}-\bm{b}-\bm{z}\|^{2}$
$\qquad-\frac{1}{2}\|\bm{z}\|^{2}-\langle\bm{z},\,\bm{b}\rangle,$
where $\bm{z}\in\mathbb{R}^{m}$ is the dual variable and $\bm{v}_{k}(\bm{z}):=\gamma_{k}^{-1}\bm{\xi}^{k}+\bm{x}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}$. Then, the dual problem of (5.3) is give by
$\max\limits_{\bm{z}\in\mathbb{R}^{m}}\left\{\,\min_{\bm{x}\in\mathbb{R}^{n},\,\bm{y}\in\mathbb{R}^{m}}~{}\mathcal{L}(\bm{x},\bm{y},\bm{z})\,\right\}.$
Observe that
$~{}\min_{\bm{x}\in\mathbb{R}^{n},\,\bm{y}\in\mathbb{R}^{m}}~{}\mathcal{L}(\bm{x},\bm{y},\bm{z})$
$=\min_{\bm{x}\in\mathbb{R}^{n}}\left\{\lambda\|\bm{x}\|_{1}+\frac{\gamma_{k}}{2}\|\bm{x}-\bm{v}_{k}(\bm{z})\|^{2}\right\}-\frac{\gamma_{k}}{2}\left\|\bm{v}_{k}(\bm{z})\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{x}^{k}\|^{2}$
$\qquad+\min_{\bm{y}\in\mathbb{R}^{m}}\left\{\frac{1}{2}\|\bm{y}-\bm{b}-\bm{z}\|^{2}\right\}-\frac{1}{2}\|\bm{z}\|^{2}-\langle\bm{z},\,\bm{b}\rangle$
$=\lambda\left\|\texttt{prox}_{\lambda\gamma_{k}^{-1}\|\cdot\|_{1}}\!\left(\bm{v}_{k}(\bm{z})\right)\right\|_{1}+\frac{\gamma_{k}}{2}\left\|\texttt{prox}_{\lambda\gamma_{k}^{-1}\|\cdot\|_{1}}\!\left(\bm{v}_{k}(\bm{z})\right)-\bm{v}_{k}(\bm{z})\right\|^{2}$
$\qquad-\frac{\gamma_{k}}{2}\left\|\bm{v}_{k}(\bm{z})\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{x}^{k}\|^{2}-\frac{1}{2}\|\bm{z}\|^{2}-\langle\bm{z},\,\bm{b}\rangle.$
Thus, the dual problem of (5.3) admits the following form
$\max\limits_{\bm{z}\in\mathbb{R}^{m}}\!\left\{\begin{aligned} &-\frac{1}{2}\|\bm{z}\|^{2}-\langle\bm{z},\,\bm{b}\rangle+\lambda\left\|\texttt{prox}_{\lambda\gamma_{k}^{-1}\|\cdot\|_{1}}\!\left(\bm{v}_{k}(\bm{z})\right)\right\|_{1}\\[3.0pt] &~{}~{}+\frac{\gamma_{k}}{2}\left\|\texttt{prox}_{\lambda\gamma_{k}^{-1}\|\cdot\|_{1}}\!\left(\bm{v}_{k}(\bm{z})\right)-\bm{v}_{k}(\bm{z})\right\|^{2}-\frac{\gamma_{k}}{2}\left\|\bm{v}_{k}(\bm{z})\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{x}^{k}\|^{2}\end{aligned}\right\},$
which is equivalent to (5.4), when expressed in a minimization form.
A.2 Derivation of the dual problem (5.11)
For notational simplicity, let $g(\cdot):=\|\cdot\|_{1}+\iota_{M}(\cdot)$, $\bm{s}^{k}:=\bm{x}^{k}+\gamma_{k}^{-1}\bm{\xi}^{k}$ and $\bm{b}^{k}:=A\bm{x}^{k}-\bm{b}$. Then, the associated Lagrangian function of (5.10) is given by
$\mathcal{L}(\bm{x},\bm{y},\bm{z})$ $=g(\bm{x})+\frac{\gamma_{k}}{2}\left\|\bm{x}-\bm{s}^{k}\right\|^{2}+\iota_{\kappa}(\bm{y})+\frac{\gamma_{k}}{2}\left\|\bm{y}-\bm{b}^{k}\right\|^{2}+\langle\bm{z},\bm{\bm{A}}\bm{x}-\bm{y}-\bm{b}\rangle$
$=g(\bm{x})+\frac{\gamma_{k}}{2}\left\|\bm{x}-\big{(}\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\big{)}\right\|^{2}-\frac{\gamma_{k}}{2}\left\|\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{s}^{k}\|^{2}$
$\qquad+\iota_{\kappa}(\bm{y})+\frac{\gamma_{k}}{2}\left\|\bm{y}-\big{(}\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\big{)}\right\|^{2}-\frac{\gamma_{k}}{2}\left\|\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{b}^{k}\|^{2}-\langle\bm{z},\,\bm{b}\rangle,$
where $\bm{z}\in\mathbb{R}^{m}$ is the dual variable. Then, the dual problem of (5.10) is give by
$\max\limits_{\bm{z}\in\mathbb{R}^{m}}\left\{\,\min_{\bm{x}\in\mathbb{R}^{n},\,\bm{y}\in\mathbb{R}^{m}}~{}\mathcal{L}(\bm{x},\bm{y},\bm{z})\,\right\}.$
Observe that
$~{}\min_{\bm{x}\in\mathbb{R}^{n},\,\bm{y}\in\mathbb{R}^{m}}~{}\mathcal{L}(\bm{x},\bm{y},\bm{z})$
$=\min_{\bm{x}\in\mathbb{R}^{n}}\left\{g(\bm{x})+\frac{\gamma_{k}}{2}\left\|\bm{x}-\big{(}\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\big{)}\right\|^{2}\right\}-\frac{\gamma_{k}}{2}\left\|\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{s}^{k}\|^{2}$
$\qquad+\min_{\bm{y}\in\mathbb{R}^{m}}\left\{\iota_{\kappa}(\bm{y})+\frac{\gamma_{k}}{2}\left\|\bm{y}-\big{(}\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\big{)}\right\|^{2}\right\}-\frac{\gamma_{k}}{2}\left\|\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{b}^{k}\|^{2}-\langle\bm{z},\,\bm{b}\rangle$
$=\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|_{1}+\frac{\gamma_{k}}{2}\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)-\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|^{2}$
$\quad-\frac{\gamma_{k}}{2}\left\|\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{s}^{k}\|^{2}$
$\quad+\frac{\gamma_{k}}{2}\left\|\Pi_{\kappa}\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)-\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)\right\|^{2}-\frac{\gamma_{k}}{2}\left\|\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right\|^{2}+\frac{\gamma_{k}}{2}\|\bm{b}^{k}\|^{2}-\langle\bm{z},\,\bm{b}\rangle,$
where $\texttt{prox}_{\gamma_{k}^{-1}g}$ is the proximal mapping of $\gamma_{k}^{-1}g$ and $\Pi_{\kappa}$ is the projection operator over $\{\bm{x}\in\mathbb{R}^{n}:\|\bm{x}\|\leq\kappa\}$. Thus, the dual problem of (5.10) admits the following form
$\max_{\bm{z}\in\mathbb{R}^{m}}\left\{\begin{aligned} &-\langle\bm{z},\bm{b}\rangle-\frac{\gamma_{k}}{2}\left\|\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right\|^{2}+\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|_{1}\\[3.0pt] &\qquad+\frac{\gamma_{k}}{2}\left\|\texttt{prox}_{\gamma_{k}^{-1}g}\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)-\left(\bm{s}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}\right)\right\|^{2}\\[3.0pt] &\qquad-\frac{\gamma_{k}}{2}\left\|\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right\|^{2}+\frac{\gamma_{k}}{2}\left\|\Pi_{\kappa}\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)-\left(\bm{b}^{k}+\gamma_{k}^{-1}\bm{z}\right)\right\|^{2}\\[3.0pt] &\qquad+\frac{\gamma_{k}}{2}\|\bm{s}^{k}\|^{2}+\frac{\gamma_{k}}{2}\|\bm{b}^{k}\|^{2}\end{aligned}\right\},$
which is equivalent to (5.11), when expressed in a minimization form.
A.3 Proof of Proposition 5.1